### v1

# CyberHost: Taming Audio-driven Avatar Diffusion Model with Region Codebook Attention

Abstract
Diffusion-based video generation technology has advanced significantly, catalyzing a proliferation of research in human animation. However, the majority of these studies are confined to same-modality driving settings, with cross-modality human body animation remaining relatively underexplored. In this paper, we introduce CyberHost, an end-to-end audio-driven human animation framework that ensures hand integrity, identity consistency, and natural motion. The key design of CyberHost is the Region Codebook Attention mechanism, which improves the generation quality of facial and hand animations by integrating fine-grained local features with learned motion pattern priors. Furthermore, we have developed a suite of human-prior-guided training strategies, including body movement map, hand clarity score, pose-aligned reference feature, and local enhancement supervision, to improve synthesis results. To our knowledge, CyberHost is the first end-to-end audio-driven human diffusion model capable of facilitating zero-shot video generation within the scope of human body. Extensive experiments demonstrate that CyberHost surpasses previous works in both quantitative and qualitative aspects.
1 Introduction
Human animation aims to generate realistic and natural human videos from a single image and control signals such as audio, text, and pose sequences. In audio-driven settings, previous works (Prajwal et al., 2020; Yin et al., 2022; Wang et al., 2021; Ma et al., 2023; Zhang et al., 2023; Chen et al., 2024; Xu et al., 2024c; b; Tian et al., 2024; Wang et al., 2024a) have primarily focused on generating portrait videos from various modalities, often overlooking the challenges associated with animating the human body below the shoulders. Recently, advancements in diffusion models have led some studies (Karras et al., 2023; Wang et al., 2024b; Hu et al., 2023; Zhang et al., 2024; Xu et al., 2023; Huang et al., 2024; Corona et al., 2024) to explore their potential for enhancing full-body human video generation. However, these diffusion models are predominantly tailored for video-driven settings and do not seamlessly translate to audio-driven scenarios. In the realm of portrait animation, works like EMO (Tian et al., 2024) have demonstrated that an end-to-end audio-driven diffusion model can generate highly expressive results, yet this approach remains unexplored for full-body animation. This paper aims to address this gap.
Compared to portrait, the challenge of audio-driven body animation primarily lies in two aspects: (1) Critical human body parts such as the face and hands occupy only a small portion of the frame, yet they carry the majority of the identity information and semantic expression. Unfortunately, neural networks often fail to spontaneously prioritize learning in these key regions, making them more prone to artifacts.
(2) The correlation between audio signals and body animation control is relatively weak. Relying solely on audio signals can lead to significant uncertainty in motion generation, thereby exacerbating instability in the generated results. As shown in Figure 1, even current state-of-the-art video-driven methods still struggle with issues such as face and hand region synthesis. This challenge becomes even more severe in audio-driven setting due to the weak correlation between audio signals and motion.
Figure 1: Existing body animation methods struggle to generate detailed hand and facial results. In contrast, our approach ensures hand integrity and facial identity consistency.
For the first challenge, most existing works (Zhang et al., 2024; Huang et al., 2024) focus on enhancing the ability of diffusion models to generate key local regions, such as hands and faces, in a video-driven setting. They utilize skeleton priors as conditional inputs to simplify the modeling of body shape and motion patterns, allowing the model to focus primarily on reconstructing body textures. However, video-driven body animation typically requires motion generation modules to create pose templates, and employ retargeting technologies to address skeletal discrepancies, making it inconvenient to use. Additionally, most optimization techniques tailored for video-driven settings cannot be seamlessly adapted to audio-driven settings. Regarding the second challenge, a few studies (Liao et al., 2020; Wang et al., 2023b; Hogue et al., 2024; Corona et al., 2024) have attempt to reduce uncertainty by explicitly establishing the connection between audio signals and motion through a two-stage system. It consists of an audio-to-pose module and a pose-to-video module, using poses or meshes as intermediate representations. Nevertheless, this approach faces several critical limitations: (1) The two-stage framework design increases system complexity and reduces the model’s learning efficiency. (2) The poses or meshes carry limited information related to expressiveness, constraining the model’s ability to capture subtle human nuances. (3) Potential inaccuracies in pose or mesh annotations can diminish the model’s performance. Therefore, there is an urgent need to explore how to optimize the generation quality of key human regions within a one-stage audio-driven framework, while also reducing instability issues caused by the motion uncertainty.
In this work, we build an one-stage audio-driven body animation framework capable of zero-shot human videos generation. The framework incorporates the Region Codebook Attention mechanism to enhance the generation quality of key human regions, namely the hands and face. Specifically, the Region Codebook Attention employs a learnable spatio-temporal memory bank as motion codebook, which is guided by a learned region mask to capture common human local details from the data, including topological structures and motion patterns. Additionally, it integrates appearance features from local cropped images, serving as identity descriptors, thereby constructing a local representation that balances general details with identity-specific details for each human region. Furthermore, to address the weak correlation between audio and body motion, we designed a suite of human-prior-guided training strategies. For the inputs, we introduce the body movement map and hand clarity score as control condition to indicate body movements and hand motion states, respectively. We also utilize the skeleton map of the reference image to extract pose-aligned reference features, which indicate the current pose state of the reference image. In terms of loss functions, we designed auxiliary keypoint loss and local reweight loss for region supervision to enhance the synthesis results of local regions. In our experiments, we validated the effectiveness of the Region Codebook Attention mechanism. Combined with the proposed training strategies, CyberHost achieves superior results compared to existing methods. Moreover,we validated the exceptional performance of CyberHost in various settings, including audio-driven, video-driven, and multimodal-driven scenarios, as well as its zero-shot video generation capability for open-set test images.
We summarize our technical contributions as follows: 1) We propose the first one-stage audio-driven body animation framework enabling zero-shot human body animation without relying intermediate representations such as pose sequences. 2) We crafted a Region Codebook Attention to enhance the generation quality of key local regions such as hands and faces, by including a motion codebook to learn local structural priors and an identity descriptor to supplement appearance-related features. 3) A suite of human-prior-guided training strategies is proposed to optimize the training of human video generation in audio-driven scenarios.
2 Related Work
Video Generation. Benefiting from the advancements in diffusion models, video generation has made significant progress. Some works (Singer et al., 2022; Blattmann et al., 2023a; Zhou et al., 2022; He et al., 2022; Wang et al., 2023a) have extended the 2D U-Net pretrained on image-to-text task to 3D, achieving temporal video generation. AnimateDiff (Guo et al., 2023) trained a pluggable temporal module on large-scale video data, which can be integrated into other text-to-image diffusion models fine-tuned on specific datasets to achieve temporally smooth video generation. For controllability, VideoComposer (Wang et al., 2024c) trained a composer fusion encoder used to combine flexible conditions. Compared with UNet-based methods, DiT-based approaches have shifted the model architecture from convolutional networks to transformers. Some works, such as EasyAnimate (Xu et al., 2024a), CogVideX (Yang et al., 2024), and Sora (Tim et al., 2024), have expanded the 2D DiT framework to 3D for video generation by incorporating a specialized motion module block. Additionally, to efficiently model video data, some works have designed 3D Variational Autoencoders (VAEs) (Kingma & Welling, 2013) to compress videos along both spatial and temporal dimensions. To accelerate video generation, some methods (Podell et al., 2023; Ren et al., 2024) have proposed lightweight distillation schemes, achieving real-time performance in specific domains.
Body Animation. Existing body animation approaches mainly  (Hu et al., 2023; Wang et al., 2024b; Xu et al., 2023; Karras et al., 2023; Zhou et al., 2022) focus on video-driven settings, where the driving signal is a video sequence. DreamPose (Karras et al., 2023) uses DensePose (Güler et al., 2018) as a control signal and trains a diffusion model to perform pose transfer for any given image, thereby generating human video frames sequentially. MagicAnimate (Xu et al., 2023) extends a 2D U-Net to 3D and fine-tunes it on human body data, thereby enhancing the temporal smoothness of human video generation. AnimateAnyone (Hu et al., 2023) uses skeleton maps as control signals for its diffusion model and employs a dual U-Net architecture to maintain consistency between the generated video and the reference images. Some speech-driven body animation works, GAN-based or Diffusion-based  (Liao et al., 2020; Ginosar et al., 2019; Wang et al., 2023b; Corona et al., 2024) do exist, but these methods are typically two-stage processes. Speech2Gesture (Ginosar et al., 2019) first predicts body gesture sequence and then utilizes a pre-trained GAN to render it to final video. Vlogger (Corona et al., 2024) proposed two diffusion models for human-to-3d-motion and motion-to-2d-video. Two-stage methods require explicit representations as intermediate variables. While the use of intermediate representations can reduce the complexity of training, these representations are often not detailed enough, and the introduced errors lower the performance ceiling. Unfortunately, end-to-end approaches have not been widely explored so far.
3 Method
This section begins by introducing the fundamentals of diffusion models and outlining the overall structure of our proposed CyberHost framework in Section 3.1. Next, in Section 3.2, we detail the key designs of Region Code Attention, and explain how it is applied to the hand and face regions. Following this, in Section 3.3, we present our proposed training strategies aimed at enhancing the quality of generated videos in half-body conversational scenes.
3.1 Preliminary
We develop our algorithm based on the Latent Diffusion Model (LDM) (Blattmann et al., 2023b), which utilizes a Variational Autoencoder (VAE) Encoder (Kingma & Welling, 2013) $\mathcal{E}$ to transform the image $I$ from pixel space into a more compact latent space, represented as $z_{0}=\mathcal{E}(I)$. This transformation significantly reduces the computational load. During training, random noise is iteratively added to $z_{0}$ at various timesteps $t\in[1,...,T]$, ensuring that $z_{T}\sim\mathcal{N}(0,1)$. The training objective of LDM is to predict the added noise at every timestep $t$:
$L=\mathbb{E}_{z_{t},t,c,\epsilon\sim\mathcal{N}(0,1)}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c)\rVert_{2}^{2}\right],$ (1)
where $\epsilon_{\theta}$ denotes the trainable components such as the Denoising U-Net, and $c$ represents the conditional inputs like audio or text. During inference, the trained model is used to iteratively remove noise from a noised latent sampled from a Gaussian distribution. Subsequently, the denoised latent is decoded into an image using the VAE Decoder $\mathcal{D}$.
The overall architecture of our proposed CyperHost is illustrated in Figure 2. We referenced the design of the reference net from AnimateAnyone(Hu et al., 2023) and TryOnDiffusion(Zhu et al., 2023b), as well as the motion frames from Diffused-Heads(stypuflkowski2024diffused) and EMO (Tian et al., 2024), to construct a baseline framework. Specifically, a copy of 2D U-Net is utilized as reference net to extract the reference features from the reference image and motion features from the motion frames. The reference features and motion feature are injected into the Denoise U-Net through cross attention in the spatial and temporal dimensions, respectively. We extend the 2D Denoising U-Net to 3D by integrating the pretrained temporal module from AnimateDiff (Guo et al., 2023), enabling it to predict human body video clips. Multi-stage audio features extracted by Wav2vec (Schneider et al., 2019) $\mathcal{W}$ are integrated through cross-attention to facilitate audio-driven setting. Based on the baseline, to enhance the modeling capability for the key human region, i.e., hands and faces, we adapt the proposed Region Codebook Attention (detailed in section 3.2) to both the facial and hand regions and insert them into multiple stages of the Denoising U-Net. The Region Codebook consists of two parts: the motion codebook learned from the dataset and the identity descriptor extracted from cropped local images. To reduce the uncertainty in full-body animation driven solely by audio, several improvements (detailed in section 3.3) have been implemented: (1) The Body Movement Map is employed to stabilize the root movements of the body. It is encoded and merged with the noised latent, serving as the input for the denoising U-Net. (2) Hand clarity is explicitly enhanced by incorporating the Hand Clarity Score as a residual into the time embedding to mitigate the effects of motion blur in the data. (3) The Pose Encoder encodes the reference skeleton map, which is then integrated into the reference latent, yielding a Pose-aligned Reference Feature.
Figure 2: The overall structure of CyberHost. We aim to generate a video clip by driving a reference image based on an audio signal. A Reference Net is used to extract the pose-aligned appearance features from reference image and motion cues from motion frames. Region Codebook Attention are inserted at multiple stages of the Denoising U-Net for fine-grained modeling of local regions. The Body Movement Map is used to control the motion range of the body’s root nodes, while the Hand Clarity Score is used to control the clarity of the generated hand regions.
3.2 Region Codebook Attention
As shown in Figure 3, our proposed Region Codebook Attention employs a spatio-temporal memory bank to learn motion codebook and injects identity descriptors extracted from cropped local images. The former aims to learn identity-agnostic features, while the latter focuses on extracting identity-specific features. This design can be utilized to enhance synthesis results across any region of the human body. In subsequent sections, we specifically apply it to the face and hands, two areas that present significant challenges, and have confirmed its effectiveness.
Motion Codebook. While the popular dual U-Net architecture effectively maintains overall visual consistency between the generated video and reference images, it struggles with generating fine-grained texture details and complex motion patterns in local areas like the face and hands. This challenge is further exacerbated in the task of audio-driven human body animation due to the absence of explicit control signals. To address this, we introduce a spatio-temporal memory bank to learn shared local structural priors, including common texture features, topological structures, and motion patterns. We refer to this learned spatio-temporal memory bank as motion codebook and leverage its learned structural priors to prevent local motion degradation. The motion codebook is composed of two sets of learnable basis vectors: $\mathbf{C}_{\text{spa}}\in\mathbb{R}^{1\times n\times d}$ for spatial features and $\mathbf{C}_{\text{temp}}\in\mathbb{R}^{1\times m\times d}$ for temporal features, where $n$ and $m$ denote the number of basis vectors and $c$ denotes the channel dimension. We consider the combination of $\mathbf{C}_{\text{spa}}$ and $\mathbf{C}_{\text{temp}}$ as a pseudo 3D memory bank, endowing it with the capability to learn spatio-temporal features jointly. This capability facilitates the modeling of 3D characteristics such as hand motion. Furthermore, we constrain the basis vectors of the memory bank to be mutually orthogonal to maximize the learning capacity of the motion codebook. Specifically, the Gram-Schmidt process is applied to these vectors during each forward pass.
The regional motion codebook is integrated into the U-Net through a spatio-temporal cross-attention as shown in Figure 3. Given the backbone feature $\mathbf{F}_{\text{unet}}^{\text{in}}$ from U-Net, we apply cross attention with $\mathbf{C}_{\text{spa}}$ in the spatial dimension and with $\mathbf{C}_{\text{temp}}$ in the temporal dimension. The final output $\mathbf{F}_{\text{motion}}$ is formulated as the sum of two attentions’ result,
$\mathbf{F}_{\text{motion}}$ $=\text{Attn}(\mathbf{F}_{\text{unet}}^{\text{in}},\mathbf{C}_{\text{spa}},\mathbf{C}_{\text{spa}})+\text{Attn}(\mathbf{F}_{\text{unet}}^{\text{in}},\mathbf{C}_{\text{temp}},\mathbf{C}_{\text{temp}})$ (2)
$=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}_{\text{spa}}^{T}}{\sqrt{d}}\right)\cdot\mathbf{V}_{\text{spa}}+\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}_{\text{temp}}^{T}}{\sqrt{d}}\right)\cdot\mathbf{V}_{\text{temp}}$ (3)
where $\text{Attn}(*,*,*)$ denotes the cross attention, $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are the query, key, and value, respectively, projected from the input. We aim for $\mathbf{F}_{\text{motion}}$ to fully utilize the spatio-temporal motion priors of the local region learned within the 3D memory bank, refining and guiding the U-Net features through residual addition. To effectively focus the memory bank on feature learning for the target local region while filtering out gradient information from unrelated areas, we require a regional mask to weight the residual addition process. To achieve this, and to avoid introducing additional regional mask as input, we employ auxiliary convolutional layers to directly predict a regional attention mask $\mathbf{M}_{\text{r}}$ using U-Net feature $\mathbf{F}_{\text{unet}}^{\text{in}}$.
Identity Descriptor. It is worth noting that the process of learning motion codebook is identity-agnostic. It relies on the statistical analysis of regional common structure and motion patterns within dataset. However, identity-specific features such as hand size, skin color or textures may be overlooked. To addressing this, we employ a Regional Image Encoder $\mathcal{R}$ to extract identity-aware regional features from the cropped region image $\mathbf{I}_{\text{r}}$. The extracted feature is referred to as the identity descriptor. For clarity, we illustrate this process in the left bottom of Figure3 using the hand as an example. Combining the identity-independent motion codebook and the identity descriptor, the mathematical formulation of the overall region codebook attention can be expressed as follows:
$\mathbf{F}_{\text{id}}=\text{Attn}(\mathbf{F}_{\text{unet}}^{\text{in}},\mathcal{R}(\mathbf{I}_{\text{r}}),\mathcal{R}(\mathbf{I}_{\text{r}}))$ (4)
$\mathbf{F}_{\text{unet}}^{\text{out}}=(\mathbf{F}_{\text{motion}}+\mathbf{F}_{\text{id}})*\mathbf{M}_{\text{r}}+\mathbf{F}_{\text{unet}}^{\text{in}}$ (5)
Application to Hand and Facial Region. Note that both hand and facial features can be divided into identity-independent common structural features and identity-dependent appearance features. Therefore, the design principle of region codebook attention ensures its applicability to feature modeling for both hand and facial regions. In terms of implementation details, we use a structure similar to the Pose Encoder but with deeper blocks for the Hand Encoder to enhance its feature extraction capability. The images of both hands are individually cropped and resized to a resolution of 128 for feature extraction, and then concatenated to form the hand identity descriptor. For the facial region, we utilize a pre-trained ArcFace (Deng et al., 2019) network for feature extraction. Considering the rich details of the face, facial images are resized to a resolution of 256 for feature extraction. Both hand and facial detect boxes are determined by the minimal enclosing area defined by their respective key points. During training, the hand and facial detect boxes are also utilized to supervise the learning of $\mathbf{M}_{\text{r}}$.
Figure 3: An illustration of Region Codebook Attention, using the hand region as an example. It consists of a learnable motion codebook and an identity descriptor extracted from cropped images. The motion codebook consists of a series of 2+1D memory banks ($\mathbf{C}_{\text{spa}}$ and $\mathbf{C}_{\text{temp}}$) that interact with the U-Net features($\mathbf{F}_{\text{unet}}^{\text{in}}$) through Spatio-Temporal Cross Attention. We use a Hand Encoder to extract appearance feature from the cropped hand images, serving as the hand identity descriptor. A learned hand attention map is utilized as mask to guide the focus area of hand codebook attention.
3.3 Human-Prior-Guided Training Strategy
The performance of diffusion models is closely tied to the quality of the training data. Recent Text-to-Video studies (Podell et al., 2023; Blattmann et al., 2023a) demonstrate that designing condition inputs such as resolution and cropping parameters can enhance the model’s robustness to varied data and improve the output controllability. Inspired by them, we design the Body Movement Map and Hand Clarity Score conditions to decouple hard cases in the dataset while also reducing the uncertainty caused by the weak correlation between audio and body motion. Additionally, we designed the Pose-Aligned Reference Feature and Local Enhancement Supervision to guide the model in fully considering the skeletal topology information during the video generation process.
Body Movement Map. Frequent body movements, including translations and rotations, are present in the talking body video data, which increase the training difficulty. To address this issue, we design a Body Movement Map to serve as a control signal for the movement amplitude of body root in generated videos. Specifically, we determined a rectangular box representing the motion range of the thorax point over a video segment. To avoid a strong correlation between the motion trajectory of the thorax point and the boundaries of the rectangular box, we augmented the size of the rectangular box by $100\%-150\%$. The body movement map is down-sampled and encoded through a learnable Pose Encoder and added as a residual to the noised latent. During testing, we typically input a body movement map of fixed size to ensure the stability of the overall generated results.
Hand Clarity Score. Blurry hand images tend to lose structural details, weakening the model’s ability to learn hand structures and causing the model to generate indistinct hand appearances. Therefore, we introduce a Hand Clarity Score to indicate the clarity of hand regions in the training video frames. This score is used as a conditional input to the denoising U-Net, enhancing the model’s robustness to blurry hand data during training and enabling control over the clarity of the hand images during inference. Specifically, for each frame in the training data, we crop the pixel areas of the left and right hands based on key points and resize them to a resolution of $128\times 128$. We then use the Laplacian operator to calculate the Laplacian standard deviation of the hand image frames. A higher standard deviation typically indicates clearer hand images, and we use this value as the Hand Clarity Score. The hand clarity score is provided to the U-Net model by residually adding it to the time embedding. During inference, a higher clarity score is applied to enhance the generation results for the hands.
Pose-aligned Reference Feature. Recent works (Zhu et al., 2023b; Hu et al., 2023; Tian et al., 2024) have utilized Reference Net to extract and inject appearance features from the reference image, thereby maintaining overall visual consistency between the generated video and the reference image. However, these approaches have overlooked the topological consistency of the human skeleton between the two. We propose the use of Pose-aligned Reference Feature to ensure both visual and topological consistency in the generated videos. Specifically, we achieve pose alignment by residually adding the encoded skeleton map to the reference image latent, followed by feature extraction using the Reference Net. Consequently, the extracted reference features not only include the appearance information of the human body but also incorporate its topological structure information.
Local Enhancement Supervision. To help the model better learn the intrinsic topological structure of the human body, we introduce a keypoint loss as an auxiliary supervision signal. Specifically, after each hand codebook attention, we pass the locally refined features $\mathbf{F}_{\text{unet}}^{\text{out}}$ through several convolutional layers to predict the hand keypoints heatmap $\mathbf{\hat{H}}$. Considering that the signal-to-noise ratio varies at different time steps $t$, we only apply this loss with a $50\%$ probability when timestep $t<500$. Additionally, we employed a local reweight strategy to optimize the original training objective. Considering that local regions such as the face and hands contain richer appearance details, it is crucial to focus more on the loss optimization in these areas. During training, we use keypoints to obtain mask $\mathbf{M}$ for critical regions like the face and hands and use it to reweight the training loss $L$ by a factor $\alpha$.
$L_{les}=(1+\alpha*\mathbf{M})*L+\frac{1}{N}\sum_{i=1}^{N}\lVert\mathbf{H}_{i},\mathbf{\hat{H}}_{i}\rVert_{2}^{2}$ (6)
where $\mathbf{H}$ denotes the ground truth keypoints heatmap, and $N$ denotes the number of region codebook attention modules in U-Net. We found that setting $\alpha=1$ yielded the most stable results.
4 Experiments
In this section, we first provide the implementation details of our method and the experimental setup. Following this, we conduct quantitative and qualitative comparisons with state-of-the-art methods to validate the superior performance of our approach. We also perform ablation studies to analyze the effectiveness of our modules and training strategies. Finally, we explore the effects of our method in a multimodal-driven setting and examined its generalization ability on open-set test images.
4.1 Implement details
The training process is divided into two stages. The first stage aims to teach the model how to maintain visual consistency between the generated video frames and the reference images. In this stage, two arbitrary frames from the training video clips are sampled as the reference frame and target frame, respectively. The primary training parameters include those of the Reference Net, Pose Encoder, and basic modules within the Denoising U-Net. The training was conducted for a total of 4 days on 8 A100 GPUs, with a batch size of 12 per GPU and a resolution of $640\times 384$. In the second stage, we begin end-to-end training for the task of generating videos from images and audio. During this phase, the parameters of modules such as the temporal layers, audio attention layers, and region codebook attention layers are also optimized. Each video clip has a length of 12 frames, with the motion frames’ length set to 4. We use a total of 32 A100 GPUs to train for 4 days, with each GPU processing one video sample. This setup allows us to train with different resolutions on different GPUs. We constrain these different resolutions to have an area similar to the $640\times 384$ resolution, with both the height and width being multiples of 64 to ensure compatibility with the LDM structure. Each stage is trained with the learning rate set to $1e^{-5}$. The classifier-free guidance (CFG) scale for the reference image and audio is set to 2.5 and 4.5, respectively. We used video data collected from the internet featuring half-body speech scenarios for training, amounting to a total of 200 hours and more than 10k unique identities. We designated 269 video segments from 119 identities as the test set for quantitative evaluation.
[…]
DiffGest.+MimicMo. 0.656 14.97 58.95 1515.9 0.377 0.496 13.427 0.833 23.40
CyberHost (A2V-B) 0.691 16.96 32.97 555.8 0.514 6.627 7.506 0.884 24.73
[…]
w/o Motion Codebook 0.687 16.53 37.80 643.9 0.523 6.384 7.719 0.859 21.35
w/o ID descriptor 0.690 16.95 35.83 582.9 0.422 6.418 7.669 0.881 22.64
w/o Face Codebook 0.685 16.86 35.14 612.8 0.425 6.299 7.796 0.880 24.11
w/o Hand Codebook 0.686 16.80 37.71 625.9 0.498 6.510 7.574 0.869 22.98
w/o Body Movement 0.680 16.76 39.83 668.6 0.458 6.372 7.769 0.867 27.54
w/o Hand Clarity 0.686 16.73 37.81 643.8 0.503 6.556 7.556 0.849 33.00
w/o Pose-aligned Ref. 0.683 16.66 38.32 660.0 0.487 6.498 7.684 0.870 23.18
w/o Local Enhancement 0.687 16.92 35.25 581.5 0.461 6.127 7.930 0.866 21.35
4.2 Comparisons with state-of-the-arts
Due to the limited comparable works in the Audio-Driven Talking Body setting, we modified our algorithm slightly for comparison in the Video-Driven Body Reenactment and Audio-Driven Talking Head experiment settings. This allows us to validate the advanced nature and generalizability of the CyberHost framework against some of the current state-of-the-art algorithms. For evaluation metrics, we use Fréchet Inception Distance (FID) to assess the quality of the generated video frames and Fréchet Video Distance (FVD) (Unterthiner et al., 2019) to evaluate the overall coherence of the generated videos. To assess the preservation of facial appearance, we calculate the cosine similarity (CSIM) between the facial features of the reference image and the generated video frames. We use SyncC and SyncD, as proposed in (Prajwal et al., 2020), to evaluate the synchronization quality between lip movements and audio signals in audio-driven settings. Additionally, Average Keypoint Distance (AKD) is used to measure the accuracy of actions in video-driven settings. Because the AKD cannot be used to evaluate hand quality in audio-driven scenarios, we compute the average of Hand Keypoint Confidence (HKC) as a reference metric for evaluating hand quality. Similarly, we calculated the standard deviation of hand keypoint coordinates within a video segment as the Hand Keypoint Variance (HKV) metric to represent the richness of hand movements.
Figure 4: The audio-driven taking body results of CyberHost compared to two-stage baseline.
Audio-driven Talking Body
Currently, only a few works such as Dr2 (Wang et al., 2023b), DiffTED (Hogue et al., 2024), and Vlogger (Corona et al., 2024) have adopted two-stage approaches to achieve audio-driven talking body video generation. However, these methods are not open-sourced, making it difficult to conduct direct comparisons. To better compare the effectiveness with the dual-stage method, we constructed a dual-stage audio-driven talking body baseline based on the current state-of-the-art audio2gesture and pose2video algorithms. Specifically, we trained DiffGesture (Zhu et al., 2023a) on our dataset to generate subsequent driving SMPLX (Pavlakos et al., 2019) pose sequences based on input audio and an initial SMPLX pose. Finally, the SMPLX meshes were converted into DWPose (Yang et al., 2023) key points, and MimicMotion (Zhang et al., 2024) was used for video rendering based on these key points.
As shown in Table 1, our proposed CyberHost significantly outperforms the two-stage baseline in terms of image quality, video quality, facial consistency, and lip-sync accuracy in the audio-driven talking body (A2V-B) setting. Figure 4 also presents a visual comparison between CyberHost and the two-stage baseline. Additionally, we utilized reference images and audio from 30 demos displayed on the Vlogger homepage to conduct both quantitative and qualitative comparisons with Vlogger. Notably, since most of Vlogger’s test videos exhibit minimal motion, the HKC indicator is relatively high, whereas the HKV indicator, which measures the diversity of movements, is very low, as shown in Table 1. As depicted in Figure 4, our proposed CyberHost surpasses Vlogger in both generated image quality and the naturalness of hand movements.
Video-driven Body Reenactment
We adapt our method to perform video-driven human body reenactment (V2V-B) by utilizing DWPose (Yang et al., 2023) to extract full-body keypoints from videos and replace the body movement maps with a sequence of skeleton maps. As shown in Table 2, we compared our CyberHost with several state-of-the-art zero-shot human body reenactment methods, including Disco (Wang et al., 2024b), AnimateAnyone (Hu et al., 2023) and MimicMotion (Zhang et al., 2024). CyberHost significantly outperforms the current state-of-the-art methods in various metrics such as FID, FVD, and AKD. The visual results in Figure 5 also demonstrate that CyberHost achieves better structural integrity and identity consistency in local regions such as the hands and face compared to current state-of-the-art results.
Table 2: Quantitative comparison with existing video-driven body reenactment methods.
[…]
Disco (Wang et al., 2024b) 0.660 17.33 57.12 1490.4 0.227 9.313
AnimateAnyone (Hu et al., 2023) 0.737 20.52 26.87 834.6 0.347 5.747
MimicMotion (Zhang et al., 2024) 0.684 17.96 23.43 420.6 0.340 8.536
CyberHost (V2V-B) 0.782 21.31 20.04 181.6 0.458 3.123
Figure 5: Comparisons with other video-driven body reenactment results
Audio-driven Talking Head.
Although our framework is designed for human body driving, it requires only minor modifications to be adapted for audio-driven talking head (A2V-H) setting. We removed the unnecessary Hand Codebook Attention and adjusted the cropping area of the training data to focus around the face. As shown in Table 3, we compared our method with Hallo (Xu et al., 2024b), VExpress (Wang et al., 2024a) and EchoMimic (Chen et al., 2024). We randomly sampled 100 videos from CelebV-HQ (Zhu et al., 2022) as the test set. Experimental results demonstrate that CyberHost achieves or surpasses current state-of-the-art performance across multiple metrics, including FID, FVD, CSIM, and Sync score.
4.3 Ablation Study
Analysis of Region Codebook Attention. As shown in Table 1, we conducted ablation experiments to analyze the structure and effectiveness of Region Codebook Attention. As we can see, the motion codebook significantly improves metrics related to image quality, such as FID and FVD, as well as metrics related to the quality of hand generation, such as HKC. The identity descriptor, on the other hand, is more closely associated with the CSIM metric, demonstrating its effectiveness in maintaining identity consistency. Additionally, we separately investigated the overall effectiveness of face codebook attention and hand codebook attention. Face codebook attention significantly improves facial-related metrics such as CSIM and Sync score. Hand codebook attention effectively reduces artifacts caused by hands, thereby enhancing image quality metrics such as FVD and FID.
Analysis of Human-prior-guided Training Strategies. We also validated the effectiveness of various human-prior-guided training strategies in Table 1. The body movement map enhances the stability of the generated human body videos, leading to improved overall video quality. This subsequently boosts metrics such as SSIM, PSNR, FID, and FVD. The hand clarity score can reduce artifacts caused by rapid hand movements and enhance hand clarity, thus significantly impacting the AKC metric. The pose-aligned reference features, by leveraging the topological structure priors of the reference images, also enhance the stability of the generated results, providing significant benefits in metrics such as AKC, FID, and FVD. Local enhancement supervision improves facial consistency, lip synchronization, and hand quality. Consequently, it significantly impacts metrics such as CSIM, Sync score, and AKC.
Table 3: Comparison with existing audio-driven talking head methods.
[…]
EchoMimic (Chen et al., 2024) 0.619 17.468 35.37 828.9 0.411 3.136 10.378
VExpress (Wang et al., 2024a) 0.422 10.227 65.09 1356.5 0.573 3.547 9.415
Hallo (Xu et al., 2024b) 0.632 18.556 35.96 742.9 0.619 4.130 9.079
CyberHost(A2V-H) 0.694 19.247 25.79 552.6 0.581 4.243 8.658
Figure 6: Multimodal-driven talking body video generation.
4.4 Multimodal-driven Video Generation
Our proposed CyberHost also supports combined control signals from multiple modalities, such as 2D hand keypoints and audio. As shown in Figure 6, the hand keypoints from Hand Pose Template are used to control hand movements and audio information is used to drive head movements, facial expressions, and lip synchronization. This driving setup leverages the explicit structural information provided by hand pose templates to enhance the stability of hand generation, while significantly improving the correlation and naturalness of head movements, facial expressions, and lip synchronization with the audio.
4.5 Audio-driven Results in Open-set Domain
To validate the robustness of the CyberHost algorithm, we tested the audio-driven talking body video generation results on open-set test images. As shown in the Figure 7, our proposed method demonstrates good generalization across various characters and is capable of generating complex gestures, such as hand interactions.
Figure 7: The audio-driven taking body results of CyberHost on the open-set test images.
5 Conclusion
This paper introduces an one-stage audio-driven talking body generation framework, CyberHost, designed to produce human videos that match the input audio with high expressiveness and realism. CyberHost features an innovative Region Codebook Attention module to enhance the generation quality of key local regions, such as hands and faces. This module uses a spatio-temporal memory bank as a motion book to provide implicit guidance for maintaining coherent topological structures and natural motion patterns. Additionally, it injects appearance features from locally cropped images as identity descriptors to ensure local identity consistency. Combined with a suite of human-prior-guided training strategies suited to reduce the motion uncertainty in audio-driven setting, including the body movement map, hand clarity score, pose-aligned reference feature, and local enhancement supervision, our CyberHost algorithm can generate stable, natural, and realistic talking body videos and achieve zero-shot human image animation in open-set domain.
Acknowledgments
We extend our gratitude to Dr. Pengfei Wei for assisting in setting up the two-stage audio-driven talking body baseline.